OpenStudy (rayeanne6):

Find the roots of the polynomial equation x^3-4x^2+x+26 A. 3+2i, 3-2i, -2 B. 3+2i, 3-2i, 2 C.-3+2i, -3-2i, 2 D. -3+2i, -3-2i, -2

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle x^3-4x^2+x+26 }\) \(\large\color{#000000 }{ \displaystyle x^3-6x^2+2x^2+13x-12x+26 }\) \(\large\color{#000000 }{ \displaystyle x^3-6x^2+13x+2x^2-12x+26 }\) \(\large\color{#000000 }{ \displaystyle x(x^2-6x+13)+2(x^2-6x+13) }\)

OpenStudy (solomonzelman):

can you go from there?